# The three phases of self-gravitating scalar field ground states

### v2

@@ Introduction
The profile of a non-isolated ground state is even more complex than equation 17—in addition to $m_{j}$ and $M_{j}$, the field $\psi_{j}$ depends on the other fields $\psi_{k\neq j}$. Although the exact dependence is nontrivial, the overall trend of $r_{s}\propto m_{j}^{-2}M_{j}^{-1}$ is preserved Schive et al. (2014). When $\lambda_{ij}=0$, the only interspecies feedback is through the gravitational field, and so multi-species equilibria are essentially multiple coincident single-field solitons, scaled to account for a background gravitational field. Therefore, it makes sense that previous work on equilibria of multi-species scalar fields has assumed spherical symmetry and coincident centers of mass Gosenca et al. (2023); Glennon et al. (2023); Luu et al. (2023a); Huang et al. (2023); van Dissel et al. (2024). In keeping with previous literature, we will call such configurations nested Luu et al. (2023b).
@@
Here, we take the final step by assuming $\lambda_{ij}>0$ and finding the ground states of such systems numerically. We use a Mathematica module named nSPIRal to find numerical solutions with multiple species and self-interactions. nSPIRal is an extension of the code used in Ref. Zagorac et al. (2021) to evolve the spherically symmetric version of the Schrödinger-Poisson system in time. Unlike the previous version, nSPIRal allows for multiple axion fields, self- and inter-field interactions, and the choice of real or imaginary time. In this work, we use it to evolve the spherically symmetric equations of motion for two fields forward in imaginary time to arrive at the overall system’s ground state; see the Appendix for more details.
@@
We find this to be a reasonable approach for the following reasons. Our initial guess for each wavefunction is some superposition of eigenstates of the system. Each eigenstate has an associated eigenenergy, with the ground state having the lowest energy. If we ignore the backreaction on the gravitational potential for a moment, each wavefunction $\psi_{j}$ will evolve as
$\psi_{j}(t)=\sum_{i=1}^{N}c^{i}_{j}\exp\Biggl{(}-iE_{j}^{i}t\Biggr{)}\phi_{j}^{i}\,,$ (20)
@@
where $\phi_{j}^{i}$ is an eigenstate, $E_{j}^{i}$ its associated eigenenergy, and $c_{j}^{i}$ a complex expansion coefficient. Now, evolving in imaginary time $r\rightarrow it$ causes the exponential to decay at a rate set by $E_{j}^{i}$. Given the ground state has the largest energy, excited states will exponentially decay first, leaving us with the ground state profile.
@@
For each simulation, we must first specify initial conditions for the $\psi_{j}$ profiles; we use the multifield solitons described in Glennon et al. (2023). We have found that much of our parameter space is relatively insensitive to changes in these three parameters; however, choosing them becomes crucial when $m_{1}=m_{2}$ and $M_{1}=M_{2}$. In this case, the imaginary time evolution preserves the symmetry $\psi_{1}(r)=\psi_{2}(r)$, and only nested soliton-like solutions are found. It is possible to find solutions that violate this symmetry by starting with $\psi_{1}(r)\neq\psi_{1}(r)$. The choice of initial conditions becomes important in this region, as the ground states of the two fields become very nearly degenerate (or exactly degenerate in the case of $m_{1}=m_{2}$ and $M_{1}=M_{2}$). We thus find a set of solutions that are still nested, but now no longer have coincident maxima; we call this type of solution nested hollow, and the type described above nested solid for disambiguation. The transition from solid to hollow profiles with increasing $\lambda_{ij}$ is illustrated in Fig. 2.
@@
Figure 2: Density profiles of numerically calculated spherically symmetric ground states. The solid lines are densities of $\psi_{1}$ and the dashed lines are densities of $\psi_{2}$. The color scale indicates the interaction strength $\lambda_{12}$; large positive values are repulsive. In each case, $M_{1}=M_{2}=50\mathcal{M}$, $m_{1}=m_{0}$ and $m_{2}=10^{-0.1}m_{0}$. When the interaction strength is small, solutions are nested, similar to those presented in figure 4 and discussed elsewhere in the literature. The less massive field $\psi_{2}$ has a larger characteristic radius, as expected from equation 17. When interactions are strongly repulsive, we find hollow solutions, in which $\psi_{2}$ has a local minimum at $r=0$.
We construct a phase diagram by testing a range of parameters $\lambda_{ij}$, $m_{1}/m_{2}$, and $M_{1}/M_{2}$. In the region where the fields are nearly degenerate, we verify that we have reached the true ground state by running several choices of initial conditions and choosing the one that minimizes the energy of the system. The results are shown in Fig. 3.
@@
Having identified two miscible yet qualitatively different states, we now recall that analytic arguments presented in the last section support a third: an immiscible phase we call separate, where the centers of mass of the two scalar field solitons are separated by a distance $d$. Such a configuration breaks spherical symmetry, and is therefore beyond the reach of our 1D imaginary time solver; however, we can use energy arguments to find the transition parameters instead. Recall that the gravitational potential energy consists of contributions from the self-gravitating configurations of each scalar field and a two-body term. The two-body gravitational energy is minimized when $d=0$, while the interaction energy is minimized as $d\rightarrow\infty$. The quantum kinetic energy is agnostic with respect to the two fields’ centers of mass, but is minimized when both density profiles are solitonic; the self-gravitating terms behave similarly. Therefore, the nested solid state arises when gravitational and quantum kinetic energies are minimized at the expense of the interaction energy; the nested hollow state arises when the interaction energy is lowered at the expense of the quantum energy of one of the states (whichever is “hollow”); and the separate state arises when the interaction and quantum energies are minimized at $d>0$ at the expense of the two-body gravitational energy.
@@
$E_{KQ}+E_{\rm{int}}+\sum|E_{\rm{grav}}^{i}|\sim|E_{\rm{grav}}^{12}|.$ (21)
@@
Furthermore, we verified the hollow phase is not just a numerical artifact of the 1D imaginary time solver by putting a range of corresponding wavefunction profiles on a 3D grid using UltraDark.jl Musoke (2024). For each case, we compared the energy of our profiles with the energy of two solitonic profiles of appropriate masses at different distances. For many cases classified as “Hollow 1” or “Hollow 2” in Fig. 3, the hollow phase did indeed minimize the total energy.
Three examples of energy-minimizing configurations verified in 3D are illustrated in Fig. 1. When the separate phase was preferred, we evaluated the energy of solitons at different distances $d>0$ to find the value which minimizes the overall energy. This is how we arrived at the profile shown in the rightmost panel of Fig. 1. We also observed a general trend of the energy-minimizing distance $d$ growing with increasing $\lambda_{ij}$. However, it should be noted that the total energy is very sensitive to numerical resolution, which is necessarily smaller in the 3D simulations than their 1D counterparts. For this reason, careful construction of a full 3D phase diagram is relegated to work in our near future.
@@
Figure 3: Classification of spherically symmetric solutions found by relaxation in imaginary time, as a function of interaction strength $\Lambda_{12}$ and particle mass ratio $m_{2}/m_{1}$ (left) and total mass ratio $M_{2}/M_{1}$ (right). Points are classified as “solid” if both fields have a local maximum at $r=r_{\text{min}}$, or “hollow 1” (“hollow 2”) if $\psi_{1}$ ($\psi_{2}$) has local minimum at $r=r_{\text{min}}$. The dotted line shows the analytic prediction $\Lambda_{12}=0.09$ for the transition from solid to hollow states. The dashed line shows the expected separation between “hollow 1“ and “hollow 2“ states at $r_{s,1}=r_{s,2}$. In the left panel, all points have $M_{1}=M_{2}=50\mathcal{M}$ and $m_{1}=m_{0}$; in the right, $M_{1}=50\mathcal{M}$ and $m_{1}=m_{2}=m_{0}$.
Discussion
@@
This work is a proof of concept that in scenarios with multiple gravitating scalar fields, there are at least three distinct phases which has phenomenological implications, including an immiscible phase. Work is ongoing to calculate ground states without the assumption of spherical symmetry. These calculations will allow a phase diagram to be constructed that includes the immiscible phase. Our work presented here indicates that relaxing the assumption of spherical symmetry results in the two fields separating into axially symmetric states; but with three or more fields even axial symmetry is broken in the immiscible phase. In general the structure of ground state configurations for $n$ fields with self- and interspecies interactions can become very complicated.
@@
Work to understand the phase diagram in such models is realistic and timely, as understanding the range of possible halo profiles in the axiverse could be confronted with data from current observational efforts constraining CDM. For example, the ongoing search for dark matter halos below $10^{6}\,M_{\odot}$ is meant to test CDM’s prediction that halos are self-similar at all scales, but could be equivalently leveraged to impose limits on axiverse particle numbers and masses given a range of possible interaction strengths and, therefore, halo shapes.
Thus, in light of the increasing diversity of equilibria in scalar dark matter, we propose that “soliton” is no longer a sufficiently descriptive term for the cores predicted to exist in ULDM models. Not only can each field’s ground state differ from the solitonic profile approximated by Eq. 17 in terms of appropriate slope and radius $r_{s}$, one field in the nested hollow phase ceases resembling a core of any kind. Furthermore, there is some disagreement in the still nascent literature on whether solitons or cores even form in multi-species scenarios. Jain et al. Jain et al. (2023) find that solitonic cores condense much more slowly when there are multiple fields—the condensation of one field serving to “frustrate” the efforts of the other—while Luu et al. Luu et al. (2023b) found that condensation can happen more quickly than in the case of a single field. The addition of non-gravitational interactions could either speed or slow this process further; either would have an effect on cosmological structure formation in the axiverse, and would therefore be testable with current and upcoming observational missions.

### v3

@@ 3 Numerical results and phase diagram
Having established an expectation for a critical interspecies interaction value $\lambda_{12}^{*}$ where a phase transition will occur, we now relax our Gaussian ansatz to find more realistic ground state profiles. Previous work suggests a starting point: in the absence of interspecies interactions, multi-species equilibria are essentially multiple coincident single-field solitons, scaled to account for a background gravitational field  [37, 38, 39, 41, 67]. The overall ground states display spherical symmetry and coincident centres of mass for each field’s soliton. The same follows from our energy argument above: when $\lambda_{ij}=0$, the only interspecies feedback is through the gravitational field and is strictly attractive. Although the exact dependence of these solutions each parameter choice is nontrivial, the overall trend of $r_{s}\propto m_{j}^{-2}M_{j}^{-1}$ from the single-field case is preserved [10]. In keeping with previous literature, we will call such configurations nested [44].
@@
Let us now consider cases with $\lambda_{ij}>0$, thus introducing a competing repulsive interspecies feedback, and find the ground states of such systems numerically. We use the nSPIRal Mathematica module to find numerical solutions with multiple species and self-interactions. nSPIRal is an extension of the code used in Ref. [68] to evolve the spherically symmetric version of the Schrödinger-Poisson system in time. In this work, we use nSPIRal to evolve the spherically symmetric equations of motion for two fields forward in imaginary time to arrive at the overall system’s ground state. Unlike the previous version, our version of nSPIRal allows for multiple axion fields, self- and interfield interactions, and the choice of real or imaginary time; see Appendix B for more details.
@@
We find imaginary time evolution to be a reasonable approach if we assume there is a spectrum of eigenstates that describes the system evolved by equations 2.1 and 2.2. Any initial profile for each wavefunction is therefore some superposition of these eigenstates. Furthermore, each eigenstate has an associated eigenenergy, with the ground state having the lowest energy. Ignoring the backreaction on the gravitational potential for a moment, each wavefunction $\psi_{j}$ will evolve as
$\psi_{j}(t)=\sum_{i=1}^{N}c^{i}_{j}\exp\Biggl(-iE_{j}^{i}t\Biggr)\phi_{j}^{i}\,$, (3.1)
@@
where $\phi_{j}^{i}$ is an eigenstate, $E_{j}^{i}$ its associated eigenenergy, and $c_{j}^{i}$ a complex expansion coefficient. Now, evolving in imaginary time $t\rightarrow it$ causes the exponential to decay at a rate set by $E_{j}^{i}$. Given the ground state has the largest absolute value of its energy $|E_{0}|$, excited states will exponentially decay first, leaving us with the ground state profile. [^5] This does not evolve the system in time, so it need not be unitary. We also re-scale the mass of the system at each imaginary time step to ensure it is conserved in each field. More details of the calculation are available in the Appendix. Thus, we arrive at the ground state of the system without a priori knowing the full spectrum of our system of equations.
For each simulation, we must first specify initial conditions for the $\psi_{j}$ soliton profiles. In principle, the utility of the imaginary time solver should not keenly depend on the initial conditions, with the understanding that 1) the true ground state is present in the eigenstate makeup of the guess and 2) the numerical parameter requirements for the solver to reach the ground state, such as total evolution time, could be impacted. We initialize using the multifield soliton profiles described in [38], thus guaranteeing a large stake of the ground state at least in the $\lambda_{ij}\rightarrow 0$ limit.
@@
We have also found that much of our parameter space is relatively insensitive to changes in our numerical parameters and even exact initial profile, with one important exception: when $m_{1}=m_{2}$ and $M_{1}=M_{2}$. The imaginary time evolution preserves the symmetry $\psi_{1}(r)=\psi_{2}(r)$, and only nested soliton-like solutions are found for any value of $\lambda_{ij}$. The two fields become very nearly degenerate (or exactly degenerate in the case of $m_{1}=m_{2}$ and $M_{1}=M_{2}$). When initializing instead with some small perturbation to each profile such that $\psi_{1}(r)\neq\psi_{2}(r)$ instead, we find solutions that break this symmetry. Removing the symmetry in the initial conditions leads to a set of solutions that are still nested, but no longer have coincident maxima; rather, one has field a density maximum that is set apart from the center in a shell with a nonzero radius from the center. We call this type of solution nested hollow, and the type described above nested solid for disambiguation. The transition from solid to hollow profiles with increasing $\lambda_{ij}$ is illustrated in Fig. 3.
@@
Figure 3: Density profiles of numerically calculated spherically symmetric ground states. The solid lines are densities of $\psi_{1}$ and the dashed lines are densities of $\psi_{2}$. The color scale indicates the interaction strength $\lambda_{12}$; large positive values are repulsive. In each case, $M_{1}=M_{2}=50\mathcal{M}$, $m_{1}=m_{0}$ and $m_{2}=10^{-0.1}m_{0}$. When the interaction strength is small, solutions are nested, similar to those presented in figure 5 and discussed elsewhere in the literature. The less massive field $\psi_{2}$ has a larger characteristic radius, as expected from equation 2.10. When interactions are strongly repulsive, we find hollow solutions, in which $\psi_{2}$ has a local minimum at $r=0$.
Note that this nested hollow phase was not predicted by our analytic arguments from the previous section, which focused on changing distance between centres of mass of each field’s soliton. Here, even when keeping centers of mass coincident by construction of our numerical solver, the two fields’ density peaks become no longer co-incident. We construct a phase diagram of these two nested cases by testing a range of parameters $\lambda_{ij}$, $m_{1}/m_{2}$, and $M_{1}/M_{2}$. In the region where the fields are nearly degenerate, we verify that we have reached the true ground state by running several choices of initial conditions and choosing the one that minimizes the energy of the system. The results are shown in Fig. 4. We center our investigation on some particle mass $m_{0}$ and mass in each of the fields of $50\mathcal{M}$ for specificity and because we know this setup is well-resolved by our code. If we assume an FDM-like scenario, a particle mass of $m_{0}=10^{-22}\,\mathrm{eV/c^{2}}$ yields a mass in the field of approximately $10^{8}\,M_{\odot}$, thus also presenting an interesting parameter space.
@@
Using our imaginary time solver, we have now identified two qualitatively different spherically symmetric phases and united them under the umbrella term nested (alternatively, mixed or miscible). In addition, the analytic arguments presented in the last section support a third: an unmixed or immiscible phase we call separate, where the centers of mass of the two scalar field solitons are separated by a distance $d$. Such a configuration breaks spherical symmetry, and is therefore beyond the reach of our 1D imaginary time solver; however, we can use energy arguments to calculate the transition parameters instead. Recall that the gravitational potential energy consists of contributions from the self-gravitating configurations of each scalar field and a two-body term. The two-body gravitational energy is minimized when $d=0$, while the interaction energy is minimized as $d\rightarrow\infty$. The quantum kinetic energy is agnostic with respect to the two fields’ centers of mass, but is minimized when both density profiles are solitonic; the self-gravitating terms behave similarly. Therefore, a nested solid state arises when gravitational and quantum kinetic energies are minimized at the expense of the interaction energy; a nested hollow state arises when the interaction energy is lowered at the expense of the quantum energy of one of the states (whichever is “hollow”); and the separate state arises when the interaction and quantum energies are minimized at $d>0$ at the expense of the two-body gravitational energy.
@@
$E_{KQ}+E_{\mathrm{int}}+\sum|E_{\mathrm{grav}}^{i}|\sim|E_{\mathrm{grav}}^{12}|.$ (3.2)
@@
The existence of this separate phase raises the possibility that our nested hollow solutions are artifacts of the 1D solver used to find them. We verified that the hollow phase is not just a numerical artifact of the 1D imaginary time solver by putting a range of corresponding wavefunction profiles on a 3D grid using UltraDark.jl [69]. For each case, we compared the energy of ground state profiles generated in 1D with the energy of two solitonic profiles of appropriate masses at a range of separations, allowing us to check if there are obvious cases in which a separate’ configuration has lower energy than a ‘hollow nested’ configuration. Three examples of energy-minimizing configurations verified in 3D are illustrated in Fig. 1. At relatively large inter-species interaction strengths, there are separate configurations that have lower energy than the corresponding nested hollow configurations. This is how we arrived at the profile shown in the rightmost panel of Fig. 1. We also observed a general trend that the energy-minimizing distance $d$ grows with increasing $\lambda_{ij}$. However, at lower interaction strengths, we were did not find separate configurations that minimized the energy, suggesting that many points are correctly classified as “hollow 1” or “hollow 2” in Fig. 4.
Taken together, these findings suggest an additional region indicating a “separate” phase will appear in a version of Fig. 4 constructed with a 3D solver. However, showing this is numerically challenging. The total energy is very sensitive to numerical resolution, which is necessarily smaller in the 3D simulations than their 1D counterparts. Part of the problem lies in the competing numerical needs of each type of energy as the system transitions into the separate phase. As the two solitons of each field begin to separate, the numerical system requires very fine spatial resolution $\Delta x$ to classify when the two centres of mass cease to become co-incident at some minimum distance $d=\Delta x$. As their separation grows so does the side of the numerical box $L$ , both to accommodate $d$ and to minimize the numerical effects of the boundary of the box. However, the pressure to keep $\Delta x$ small is not lessened due to the quantum energy, which is sensitive to the gradients of each soliton’s profile. Consequently, in the parts of parameter space where we expect the hollow and separate phases to dominate the grid size required $n_{g}=\left(L/\Delta x\right)$ is very large, putting both a memory and integration time strain on the numerical calculations.
While full 3D simulations are beyond the scope of this work, we produced some preliminary 3D numerical results that support our conclusions presented above. We produced an alternate version of our phase diagram (Fig. 4) by comparing the energy of 1D solutions from nSPIRal placed on a 3D grid and the energy of two solitons placed with their centres of mass at a range of distances. The points where the separate solitons minimized the total energy for any distance $d$ greater than a single grid point was classified as “separate”, and the minimizing distance $d$ recorded. The resulting diagram maintains the hollow phase for a much reduced range of parameters and prefers the separate phase as $\lambda_{ij}$ grows. The same results also suggest the distance $d$ which minimizes the total energy similarly favors $d=0$ on when $\lambda_{ij}\rightarrow 0$ and a growing $d$ as $\lambda_{ij}\rightarrow\infty$, with some modulation from the $m_{2}/m_{1}$ and $M_{2}/M_{1}$ choice. However, the results also display features that may be the result of numerical artifacts, such as “islands” of hollow phases within the expected separate space. We expect these are due to numerical issues (discussed in the next section), but thorough verification of the results grew beyond the scope of this work. For this reason, careful construction of a full 3D phase diagram is deferred to work in our near future.
@@
Figure 4: Classification of spherically symmetric solutions found by relaxation in imaginary time, as a function of interaction strength $\Lambda_{12}$ and particle mass ratio $m_{2}/m_{1}$ (top) and total mass ratio $M_{2}/M_{1}$ (bottom). Points are classified as “solid” if both fields have a local maximum at $r=r_{\text{min}}$, or “hollow 1” (“hollow 2”) if $\psi_{1}$ ($\psi_{2}$) has local minimum at $r=r_{\text{min}}$. The dotted line shows the analytic prediction $\Lambda_{12}=0.09$ for the transition from solid to hollow states. The dashed line shows the expected separation between “hollow 1“ and “hollow 2“ states at $r_{s,1}=r_{s,2}$. In the top panel, all points have $M_{1}=M_{2}=50\mathcal{M}$ and $m_{1}=m_{0}$; in the bottom, $M_{1}=50\mathcal{M}$ and $m_{1}=m_{2}=m_{0}$.
4 Discussion
@@
This work uses 1D simulations to show that in scenarios with two gravitating scalar fields with repulsive interspecies interactions, there are at least three distinct phases of ground states with different phenomenological implications. This includes two nested or miscible phases, where the two fields’ centres of mass are coincident, and one separate or immiscible phase, where the overall system breaks spherical symmetry. Work is ongoing to calculate ground states without the assumption of spherical symmetry and use 3D simulations to further explore these scenarios. These calculations will allow a phase diagram to be constructed that includes the immiscible phase.
@@
Our results indicate that relaxing the assumption of spherical symmetry results in the two fields separating into axially symmetric states; with three or more fields, even axial symmetry will be broken in the immiscible phase. In general, the structure of ground state configurations for $n$ fields with self- and interspecies interactions can become very complicated. Naturally, this means the richness and sheer size of the available axiverse parameter space will limit the reach of our approach. But work to understand the phase diagram in cases with low species number nn—the parameter space within our reach—is realistic and timely, as the range of possible halo profiles in the axiverse could be confronted with data from current observational efforts constraining CDM. For example, the ongoing search for dark matter halos below $10^{6}\,M_{\odot}$ is meant to test CDM’s prediction that halos are self-similar at all scales, but could be equivalently leveraged to impose limits on axiverse particle numbers and masses given a range of possible interaction strengths and, therefore, halo shapes. Furthermore, independent modeling suggests that nonlinear resonance can arise between two axion species of similar masses and equilibrate their abundances, thus making the type of scenario we examine here more likely. [70]
Thus, in light of the increasing diversity of equilibria in scalar dark matter, we propose that “soliton” is no longer a universally descriptive term for the cores predicted to exist in ULDM models. Not only can each field’s ground state differ from the solitonic profile approximated by Eq. 2.10 in terms of appropriate slope and radius $r_{s}$, one field in the nested hollow phase ceases to resemble a core of any kind. At the same time, the overall density of the system most resembles a core in the nested phase, as illustrated by the middle panel in Fig. 1. Indeed, in cases such as the one considered here where the profile and shape of each wavefunction can differ significantly from the overall density, it becomes especially important to be precise which of the two quantities terms such as “soliton” or “core” refer to.
A similar disambiguation could be worthwhile when going further to consider structure formation in such a cosmology between the condensation of solitons and the formation of halo cores. One the topic of the former, Jain et al. [42] found that solitonic cores condense much more slowly when there are multiple fields—the condensation of one field serving to “frustrate” the efforts of the other—and depends keenly on the relative relative number densities and, in some cases, particle mass ratio. Their results match overlapping cases from Chen et al. [71], who considered condensation in vector dark matter. This might lead us to expect that core formation in halos proceeds more slowly in multifield scenarios due to the increased condensation time. However, Luu et al. [44] investigated non-linear structure formation in two-field cosmological simulations and found that it can proceed faster than in the one-field case when driven by the higher small-scale power in the two-field scenario. Contrasting these seemingly opposite findings yields a few guidelines for the future: firstly, that the different parameter regimes can provide different qualitative behavior (as is also argued in this paper) and, secondly, that future work may benefit from delineating soliton condensation from core formation when making predictions about cosmological structure formation. For either process, non-gravitational interactions could either speed up or slow down timescales, with either having an effect on cosmological structure formation in the axiverse or other ultralight dark sectors. Sufficient impact would be testable with current and upcoming observational missions.
There are several possible directions for future work. The most obvious next step is a better understanding of the exact nature of the ground states in the separate or immiscible phase. While we have shown that this phase exists and are confident that its ground states would exhibit axial symmetry, we have not calculated the precise solutions nor presented a full phase diagram. It is worth emphasizing that, while we have shown that there are parameters for which the spherically symmetric ground state is hollow and verified this in 3D for our range of solutions, we have not shown that they are the configurations with lowest possible energy. Achieving the numerical resolution necessary to extend our 1D imaginary time solver onto a fully realized 3D grid would be a reasonable next step towards this goal. An ideal scenario would involve comparison with completely different methods of finding the ground state (or even full spectrum) of our system; we are investigating some alternative approaches. Furthermore, models with three or more species would admit even more complex ground states. Under the assumption of spherical symmetry, there would be regimes that admit fully nested ground states, and others in which the ground state is a mixture of nested and hollow. It is also unclear how the strength of interactions required to enter an immiscible phase scales with the number of fields. For any number of species, self-interactions would additionally affect the boundaries between the miscible and immiscible phases.
Moving beyond purely stationary ground states, it is important to be aware of the gap between a mathematical ground state of the system (solitonic or otherwise) and the shape and evolution of the centre of a numerically idealized halo, let alone a cosmological one. Important effects have been demonstrated in the literature, including the soliton’s random walk motion within a halo (see e.g. [33]) and its changing shape in different backgrounds mimicking astrophysical phenomena [72, 73]. Multifield ground states embedded in halos will certainly exhibit similar modulation, whether through changing interference patters in the halo (as in e.g. [37, 38]), dependence of core stability on species ratios in individual halos ([41, 42, 43]), or other effects yet to be explored. We are particularly interested in the behavior of our nested hollow and separate phases inside halos, and plan to investigate it further. Indeed, we emphasize that this is a necessary step before confronting our stationary results with observational data collected from realistic and dynamic systems.
Finally, while the bulk of this work and its extensions considered above can be considered a mathematical phase-space exploration, the present authors’ interest in the results stems mainly from the possible astrophysical implications and links between theories of quantum gravity such as string theory, their resultant dark sectors, and observational astronomy. Thus, longer term prospects for using our results include synthesizing limits on axiverse scenarios from string theory calculations and observational constraints, as well as mapping to predictions of other soliton-forming dark sectors to unify and advance different particle theories with similar cosmological consequences.